OpenStudy (anonymous):

Which of the following is an equivalent representation of 3^-3 ? A. 1/9 B. 1/125 C. 1/625 D. 1/27

OpenStudy (wolfe8):

When a power is negative, it means the base is a denominator raised to the positive of the power. So this is the same as 1/3^3

OpenStudy (anonymous):

so the answer......

OpenStudy (shamil98):

What is 3^3, 3 x 3 x 3 ?

OpenStudy (shamil98):

if 1/3^3 = 1/3(3)(3) =

OpenStudy (shamil98):

.............................................. why

OpenStudy (calculusxy):

3^-3=0.0370370370.. When you divide 1 by 27, you get the same result.

OpenStudy (anonymous):

thx

OpenStudy (calculusxy):

No problem :)